OpenStudy (anonymous):

A farmer has 2,400 feet of fencing and wants to fence off a rectangular field that borders a straight river. He needs no fence along the river. Write the function that will produce the largest area if x is the short side of the rectangle. f(x) = 2400x − x2 f(x) = x2 − 2400 f(x) = 2x2 − 2400 f(x) = 2400x − 2x2

OpenStudy (anonymous):

I chose D

OpenStudy (anonymous):

\[\Large 2x+y=2400\] \[\Large y=2400-2x\] \[\Large A=xy\] \[\Large A=x(2400-2x)=~~2400x-2x^2\]

OpenStudy (anonymous):

Is that correct

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Is there a drawing? Without some indication, we don't know which is the short side. I'd work it both ways and see how it goes.

OpenStudy (anonymous):



OpenStudy (anonymous):

There's no drawing but that's what I presume to be the shortest side